help please to graph these eqouations by linear programming
Constraints:
12z + 8y <= 1000
1000x + 300y <= 70000
x >= 30
1 + 1.5 >= 0
profit = 1000x + 350y

Answer by stanbon(75887)

help please to graph these eqouations by linear programming
Constraints:
12z + 8y <= 1000
Shade below the line y = (-3/2)x + 125
-------------------------------------
1000x + 300y <= 70000
Shade below the line y = (-10/3)x + (70/3)
---------------------------
x >= 30
Shade to the right of x = 30
----------------------------------
1 + 1.5 >= 0
Variable is missing in your post.
------------------------------
profit = 1000x + 350y
Find the coordinates of the intersections of your
boundary lines. Substitute each x/y pair into
this objective function to determine the profit
at each point.
===============
Cheers,
Stan H.
----------------